OpenStudy (anonymous):

help me please Jose has 9 more comic books than Robin. Robin has 5 more comic books than Lee. a. If Jose has 23 comic books, how many comic books does Robin have? b. How many comic books does Lee have? A. Robin has 14 comic books; Lee has 9 comic books B. Robin has 14 comic books; Lee has 18 comic books C. Robin has 32 comic books; Lee has 37 comic books D. Robin has 9 comic books; Lee has 5 comic books

OpenStudy (anonymous):

i dont expect people to do my work for me i just need help on this on becuz im in the middle of a test

OpenStudy (anonymous):

Jose = j Robin = r Lee = l Jose has 9 more comic books than Robin. Robin has 5 more comic books than Lee. SO: j = 9 + r r = 5 + l -------------------------------- If Jose(j) has 23 comic books, how many comic books does Robin have? Solve for Robin(r). j = 9 + r 23 = 9 + r Subtract 9. r = 14 Robin has 14 comic books. -------------------------------- How many comic books does Lee have? Solve for Lee(l). You know Robin(r) has 14. SO: r = 5 + l 14 = 5 + l Subtract 5. 9 = l l = 9 Lee has 9 books.